Obsession Inc. is a publicly traded company with two business lines: perfumes and apparel. The beta of this company is 1.72. Obsession Inc. is contemplating a major restructuring in which it wants to sell off its apparel division, which has an unlevered beta of 0.71, for 20 million. It also wants to borrow an additional 25 million and buy back stock worth 45 million. After the sale of the division and the share repurchase, Obsession Inc. will have 46 million in debt and 143 million in equity outstanding. Assume a tax rate of 40.00% percent. The unlevered beta of the perfume division is

1.61

0.71

1.71

0.32

0.95
Answer #1

Given

Beta =1.72

Restructuring i.e., selling of apparel division

Sold off value = 20 million

Unlevered Beta= 0.71,

Borrowed i.e., debt = 25 million

Stock buy back = 45 million

After restructuring

Debt = 46 million

Equity = 143 million

Tax rate t= 40%

Now , first let us find out debt and equity values before restructure

Debt before restructuring = After restructure – Borrowed amount

= 46 million – 25 million = 21 million

(note: After restructure= Before debt + Borrowed amount)

Equity before restructuring = After restructure + Stock buy back

= 143 million + 45 million = 188 million

Total debt and equity before restructure = 21 million + 188 million = 209 million

Debt/Equity Ratio before restructuring = 21 million /188 million = 11.17%

Unlevered Beta before restructuring =

Beta/ [1+ {(1-t)(D/E ratio)}]

= 1.72 / (1+(1-.4)(.1117)) = 1.61

Total debt and equity after restructure =46 million + 143 million = 189 million

Let Unlevered Beta after the divisional sale be X

Unlevered beta for sold off * (sold off value/ total debt and equity value before restructure) + X * (total debt and equity value after restructure / total debt and equity value before restructure)

0.71 *(20 million /209 million) + X (189 million /209 million) = 1.61

Solving for X,

X= [1.61 – {0.71 *(20 million /209 million) }]/ (189 million /209 million)

Unlevered Beta after divisional sale = 1.705 =1.71

Therefore the Unlevered Beta of the perfume division is 1.71